### v2

@@ Abstract
Auditory scene analysis (ASA) aims to retrieve information from the acoustic environment, by carrying out three main tasks: sound source location, separation, and classification. These tasks are traditionally executed with a linear data flow, where the sound sources are first located; then, using their location, each source is separated into its own audio stream; from each of which, information is extracted that is relevant to the application scenario (audio event detection, speaker identification, emotion classification, etc.). However, running these tasks linearly increases the overall response time, while making the last tasks (separation and classification) highly sensitive to errors of the first task (location). A considerable amount of effort and computational complexity has been employed in the state-of-the-art to develop techniques that are the least error-prone possible. However, doing so gives rise to an ASA system that is non-viable in many applications that require a small computational footprint and a low response time, such as bioacoustics, hearing-aid design, search and rescue, human-robot interaction, etc. To this effect, in this work, a multi-agent approach is proposed to carry out ASA where the tasks are run in parallel, with feedback loops between them to compensate for local errors, such as: using the quality of the separation output to correct the location error; and using the classification result to reduce the localization’s sensitivity towards interferences. The result is a multi-agent auditory scene analysis (MASA) system that is robust against local errors, without a considerable increase in complexity, and with a low response time. The complete proposed MASA system is provided as a framework (publicly available at: https://github.com/balkce/masa), that uses open-source tools for sound acquisition and reproduction (JACK) and inter-agent communication (ROS2), allowing users to add their own agents.
@@ 2.4.2 New Optimization Mechanism with DOA Merger
As a reminder, the optimization task aims to modify one input parameter ($\theta_{corr}$) to optimize one observable variable ($Q$). The insight of this new optimization mechanism is as follows: at any given moment, the correction process can recognize that it has overshot the correct DOA by remembering the $\theta_{corr}$ that has provided the highest $Q$ up until to that point ($Q_{best}$), and ‘reset’ itself if necessary. Thus, if none of the recent $\theta_{corr}$ values have not provided a higher $Q$ than $Q_{best}$, the new optimization mechanism reverts back to the $\theta_{corr}$ value that provided it ($\theta_{best}$) and re-starts the optimization process from there. This new mechanism is detailed in Algorithm 1, where all variables are initialized at 0, unless stated otherwise.
@@
3:Initialize: $\theta_{hist}\leftarrow[0,0,\dots,0]$; $i_{nocormax}\leftarrow 10$; $\theta_{best}\leftarrow\mathit{None}$; $Q_{best}\leftarrow Q_{MAX}$
@@
Table 7: New optimization mechanism compared to the original.
$\theta_{error}$ $\theta_{var}$ # good runs
$\theta_{est}$ orig. new orig. new orig. new
$1^{\circ}$ $6.72$ $\mathbf{4.96}$ $3.02$ $\mathbf{2.20}$ 55 $\mathbf{6}$
$5^{\circ}$ $4.85$ $\mathbf{3.77}$ $2.51$ $\mathbf{0.54}$ 77 $\mathbf{10}$
$10^{\circ}$ $4.93$ $\mathbf{4.02}$ $2.41$ $\mathbf{0.67}$ 66 $\mathbf{9}$
$15^{\circ}$ $6.19$ $\mathbf{5.99}$ $3.63$ $\mathbf{2.67}$ 44 $\mathbf{5}$
$20^{\circ}$ $9.28$ $\mathbf{7.03}$ $6.27$ $\mathbf{3.65}$ 22 $\mathbf{4}$
$25^{\circ}$ $\mathbf{11.31}$ $12.82$ $\mathbf{7.33}$ $7.34$ 0 0
  • Reduces the error throughout the run (less $\theta_{error}$).
  • Provides considerably more consistent results (less $\theta_{var}$).
  • Increases the rate of success (more ‘good runs’).
@@ 2.5 Frequency Selection based on Source Type
In Figure 14, the output of the soundloc agent is shown when it masked the peak frequencies of the urban sound, which are [999,1229,1485,1630,2005,2478,2844,2985,3257,3486,3670,3976][999,1229,1485,1630,2005,2478,2844,2985,3257,3486,3670,3976] Hz. These were extracted from the log-frequency domain, calculated from the urban sound recording that was inserted into the input audio mixture.
@@ Acknowledgements
Declarations
[…]
Funding This work was supported by PAPIIT-UNAM through the grant IN100624.
[…]
Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
[…]
Code availability The code is available at: https://github.com/balkce/masa.
[…]

### v3

@@ Abstract
Auditory scene analysis (ASA) aims to retrieve information from the acoustic environment, by carrying out three main tasks: sound source location, separation, and classification. These tasks are traditionally executed with a linear data flow, where the sound sources are first located; then, using their location, each source is separated into its own audio stream; from each of which, information is extracted that is relevant to the application scenario (audio event detection, speaker identification, emotion classification, etc.). However, running these tasks linearly increases the overall response time, while making the last tasks (separation and classification) highly sensitive to errors of the first task (location). A considerable amount of effort and computational complexity has been employed in the state-of-the-art to develop techniques that are the least error-prone possible. However, doing so gives rise to an ASA system that is non-viable in many applications that require a small computational footprint and a low response time, such as bioacoustics, hearing-aid design, search and rescue, human-robot interaction, etc. To this effect, in this work, a multi-agent approach is proposed to carry out ASA where the tasks are run in parallel, with feedback loops between them to compensate for local errors, such as: using the quality of the separation output to correct the location error; and using the classification result to reduce the localization’s sensitivity towards interferences. The result is a multi-agent auditory scene analysis (MASA) system that is robust against local errors, without a considerable increase in complexity, and with a low response time. The complete proposed MASA system is provided as a publicly available framework that uses open-source tools for sound acquisition and reproduction (JACK) and inter-agent communication (ROS2), allowing users to add their own agents.
@@
journal: Applied Soft Computing
[…]
[1]organization=Instituto de Investigaciones en Matematicas Aplicadas y en Sistemas, Universidad Nacional Autonoma de Mexico,addressline=Circuito Escolar 3000, city=Coyoacan, postcode=04510, state=Ciudad de Mexico, country=Mexico
[…]
[2]organization=M365 Core, Microsoft Corporation,addressline=1 Microsoft Way, city=Redmond, postcode=98052, state=Washington, country=United States
[…]
[3]organization=Posgrado en Ciencias e Ingenieria de la Computacion, Universidad Nacional Autonoma de Mexico,addressline=Circuito Escolar 3000, city=Coyoacan, postcode=04510, state=Ciudad de Mexico, country=Mexico
@@ 2.4.2 New Optimization Mechanism with DOA Merger
The new optimization mechanism is such that, at any given moment, the correction process can recognize that it has overshot the correct DOA by remembering the $\theta_{corr}$ that has provided the highest $Q$ up until to that point ($Q_{best}$), and ‘reset’ itself if necessary. Thus, if none of the recent $\theta_{corr}$ values have not provided a higher $Q$ than $Q_{best}$, the new optimization mechanism reverts back to the $\theta_{corr}$ value that provided it ($\theta_{best}$) and re-starts the optimization process from there. This new mechanism is detailed in Algorithm 1, where all variables are initialized at 0, unless stated otherwise.
@@
3:Initialize: $\theta_{hist}\leftarrow[0,0,\dots,0]$; $i_{ncm}\leftarrow 10$; $\theta_{best}\leftarrow\mathit{None}$; $Q_{best}\leftarrow Q_{MAX}$
@@
$\theta_{error}$ $\theta_{var}$ # good runs
$\theta_{est}$ orig. new orig. new orig. new
$1^{\circ}$ $6.72$ $\mathbf{4.96}$ $3.02$ $\mathbf{2.20}$ 55 $\mathbf{6}$
$5^{\circ}$ $4.85$ $\mathbf{3.77}$ $2.51$ $\mathbf{0.54}$ 77 $\mathbf{10}$
$10^{\circ}$ $4.93$ $\mathbf{4.02}$ $2.41$ $\mathbf{0.67}$ 66 $\mathbf{9}$
$15^{\circ}$ $6.19$ $\mathbf{5.99}$ $3.63$ $\mathbf{2.67}$ 44 $\mathbf{5}$
$20^{\circ}$ $9.28$ $\mathbf{7.03}$ $6.27$ $\mathbf{3.65}$ 22 $\mathbf{4}$
$25^{\circ}$ $\mathbf{11.31}$ $12.82$ $\mathbf{7.33}$ $7.34$ 0 0
Table 7: New optimization mechanism compared to the original.
  • 1.
    Reduces the error throughout the run (less $\theta_{error}$).
  • 2.
    Provides considerably more consistent results (less $\theta_{var}$).
  • 3.
    Increases the rate of success (more ‘good runs’).
@@ 2.5 Frequency Selection based on Source Type
In Figure 14, the output of the soundloc agent is shown when it masked the peak frequencies of the urban sound, which are: $999$, $1229$, $1485$, $1630$, $2005$, $2478$, $2844$, $2985$, $3257$, $3486$, $3670$, and $3976$ Hz. These were extracted from the log-frequency domain, calculated from the urban sound recording that was inserted into the input audio mixture.
@@ Acknowledgements
Author Contributions
Conceptualization; Formal analysis and investigation; Writing - original draft preparation; Funding acquisition; Resources; and Supervision: Caleb Rascon. Methodology; Writing - review and editing: Caleb Rascon, Luis Miguel Gato-Díaz, Eduardo García-Alarcón.
Funding
This work was supported by PAPIIT-UNAM through the grant IN100624.